OpenStudy (anonymous):

Find two different solutions to the equation a + b i = 4 + 5 i

mathslover (mathslover):

Hello gayabee. First of all Welcome To OpenStudy.

OpenStudy (anonymous):

Why thank you :)

OpenStudy (anonymous):

i m wondering about the second solution

OpenStudy (anonymous):

Yeah, I got the first ones where a= 4 and b = 5

mathslover (mathslover):

I have : \(a + bi = 4 + 5 i \) By equating real parts of the both sides : Real Part of (a+bi) = Real part of (4+ 5i) Real part of (a+bi) = a and Real part of (4+5i) = 4 so a = 4

mathslover (mathslover):

yes and by equating imaginary parts : b = 5

OpenStudy (anonymous):

But apparently there are 2 sets of solutions that being the first set.

mathslover (mathslover):

There will only be a single set of soln .

OpenStudy (anonymous):

i also think that

OpenStudy (anonymous):

I thought that too but apparently not. I'm on my uni site doing an online quiz

mathslover (mathslover):

Do you have options?

OpenStudy (anonymous):

where a and b are real or complex numbers, and neither of a or b is zero. 1 and 2. Solution One: a = b = 3 and 4. Solution Two: a = b =

OpenStudy (anonymous):

That's all I've got

mathslover (mathslover):

Hmm, put in the first one , a = 4 and b = 5 and let the second one be empty. is it allowed to do so?

mathslover (mathslover):

Can you snap the question and attach here?

OpenStudy (primeralph):

we could extend this to another space of complex numbers. Do you understand Re and Im graphs?

OpenStudy (anonymous):

yes i do

mathslover (mathslover):

@primeralph are you sure that you can arrive to another different set of solution for a and b in a + b = 4 + 5i ? [ we have one soln as : a = 4 and b = 5 ]

mathslover (mathslover):

@gayabee please snap the quest. and post here.

OpenStudy (anonymous):

i can't seem to snap it but this is exactly what it is ----------------------------------------------------------------- Find two different solutions to the equation a + b i = 4 + 5 i where a and b are real or complex numbers, and neither of a or b is zero. 1 and 2. Solution One: a = b = 3 and 4. Solution Two: a = b =

OpenStudy (primeralph):

@mathslover simply speaking, no. Analytically speaking, yes. @gayabee perfect, but I must warn you, this isn't really an answer unless it is specially requested.

OpenStudy (anonymous):

all right. got that. thank you! @primeralph

OpenStudy (primeralph):